OpenStudy (anonymous):

solve using the multiplication principle don't forget to perform a check-3/4x=-5/6 what is the solution

jhonyy9 (jhonyy9):

-18=-20x x=-18/-20 =9/10

jhonyy9 (jhonyy9):

this is ok ?

OpenStudy (anonymous):

its not 10/9

jhonyy9 (jhonyy9):

why being 10/9 ?

jhonyy9 (jhonyy9):

we get -18=-20x yes ?

jhonyy9 (jhonyy9):

this is right ?

jhonyy9 (jhonyy9):

if yes divid both sides with -20 and you will get -18/-20 = -20x/-20 -18/.20 = -9/10 -20x/-20 =x

jhonyy9 (jhonyy9):

sorry we have -18/-20 what will be 9/10

jhonyy9 (jhonyy9):

hope that now you have understood it